Fermi National Laboratory
 

Fermilab Shutdown Report for May 5, 2006 - May 12, 2006

 

LINAC

Current LINAC Status

  • 400 MeV Beam has been established
  • H.E. Debuncher water readbacks calibrated
  • LRF5 SCR controller has been fixed
  • LRF1 Cavity water pump replaced
  • HRF3 Coil Power Supply replaced
  • LRF1 IPA1 Power Supply reset
  • Collimator installed, but motor can not pull open, needs to be operated manually
  • LRF7 (Test Stand) – work is currently ongoing to repair station and hope to return to testing 7835 tubs soon

LINAC Plans for Next Week

  • NTF Patient Treatment (M, W, & Th)
  • 400 MeV Beam for LINAC studies and Booster Studies

Finished Work

  • Cleaned L.E. LINAC Modulators & Rectifiers (LFR1-LRF5)
  • Replace water fittings in Klystron Modulator De-Qing circuit
  • Replaced Collimator between L.E. LINAC Tanks 1 & 2
  • L.E. LINAC Water Maintenance
  • Modified Chopper & Moved Quad for future MuCool operations
  • Completed McCool Alignment
  • Annual Pre-Acc Safety Trip Test
  • Cleaned Filters on 8 Ion Pump PS’s
  • 45°/90° Magnet PS Maintenance
  • New I-Anode Bulk Supply On-Line
  • BPM Digitizers Installed
  • All major L.E. LINAC LLRF system characterization data collected

Remaining Work

  • None

Booster

  • MI-8 Dump
    • All Supplies Tested
    • Polarity checks completed
    • SEPTA MP03 has a ground faulted cable – will repair today
  • 400 MeV/ORBMP
    • All supplies tested (except MV1 – safety lock)
      • MH2 – problem with some interlocks
    • ORBMP foils installed – waiting for beam tests
  • RF repair/maintenance continues – Most stations On
    • RF 6 has a bad tuner and will need to be pulled (2 days)
  • Long 13 extraction elements removed – LCW water drained
  • GMPS supplies tested - # 4 still needs some adjusting
    • Transformers Energized
  • Safety System Tests Completed

Schedule

  • Booster Beam signoff
    • New Limits on Booster BBM
  • DC Beam studies – later today/evening
  • Accelerated Beam to dump – early next week

Pbar

On Schedule

Main Injector

On Schedule

Recycler

On Schedule

Tevatron

Jobs in Process

  • F 17-3 kicker replacement
  • TEL 2 vacuum leak
  • A3 vacuum certification
  • Separator conditioning
  • TEL 1 Bake
  • A17 collimator Bake
  • E11 bake
  • HOPs cable installation
  • HLS at B0
  • HLS D0
  • A 48 collimator alignment
  • Install poly beads at B0 D0
  • Tunnel clean up

SwitchYard

Work Continues

NuMI

Work Continues

MiniBooNE

  • Work Completed
    • Horn test.
    • RAW system.
    • Triplet realigned.
    • Desiccator filter changed.
    • Belleville washers complete.
    • Leaky magnets fixed.
  • What’s Left?
    • Need signatures.
    • Safety system test.
    • Begin running power supplies.
    • Begin pulsing horn Monday (May 15).

Collider Detector at Fermilab

On schedule

On Schedule

D-Zero (D0)

  • Friday May 12
    • Upgraded Silicon Micro strip Tracker readout verification
  • Monday May 15
    • Reconfigure detector in preparation for beam pipe installation
  • Tuesday May 16
    • Install SNEG beam pipes (AD support)
  • Wednesday May 17
    • Leak check beam pipes (AD support)
  • Thursday May 18 and Friday May 19
    • Activate SNEGs (assuming leak checking was successful)
  • Friday May 19
    • Survey End Calorimeters after closing (AMG support-5169)
  • Monday May22
    • Detector checkout
  • Tuesday May 23
    • Detector checkout
  • Wednesday May 24
    • Close west CF and monitor for calorimeter noise
  • Thursday May 25
    • Close east CF and EFs
  • Friday May 26
    • Rollout nose pieces and close clam shells
    • VStar Survey detector center beam (AMG support-5318)
  • Tuesday May 30
    • VStar Survey detector center beam (AMG support-5318)
    • Search and secure collision hall during early evening for power on tests of solenoid and toroid (AD ops support)
  • Wednesday May 31
    • Closing EF, CF stick mic survey (AMG support-5317)

Cryogenics

Cryo Operation

  • Tevatron at LN2 temperature and LN2 usage is stable with vendor deliveries running only 2-3 truckloads per day due to the Tevatron being 1/6 at room temperature. Helium losses are slightly down at 12 kscf/day (1/3 of normal during TeV operations); some helium being replenished from A0 GHe Recovery of test caves.
  • A1 and A3 are at room temperature and isolated with LOTO.

D1 and F4 are also at room temperature, have been reconnected by Cryo, and are under helium flow. D1 cooldown permit signoff has been completed and cooldown to LN2 temperature will commence soon.

F4 cooldown will wait until TEL-1 is re-installed.

A3 has been completed by MSD, but spool piece conning towers retrofit by TD is in progress.

  • Pbar D60 Debuncher tanks were cooled down to LHe temperature after Division sign-off.
  • ACNET application programs relating to FRIG operations continue to be tested on Linux consoles.
  • Nitrogen Reliquefier plant was started and cooled down successfully. It was shut off later for some minor repairs and due to lack of sufficient nitrogen gas coming from Tevatron due to E0 still being on generator power. We plan to cool it down again next week.
  • CHL Helium Compr “C” was started for testing and purification but was shut down due to 2 nd stage suction cylinder valve failure.
  • MCR monitoring cryo operations but CHL 24/7 shiftwork operations will resume on Monday 5/15/2006 0600.

CHL Maintenance:

  • Helium Compr motor starter switchgear had mechanical binding that was addressed by FESS. Compr “A” was started fine thereafter.
  • Helium Compr “D” overhaul is almost completed but awaiting installation of cylinder valves that are being delivered today.
  • Helium Coldbox-2 controls upgrade is in progress.

FRIG Maintenance:

  • Various FRIG bldg vacuum maintenance and instrumentation repairs (e.g. lost charge on VPTs) are in progress, as well as support work on houses at room temperature and other equipment as things fail even during the shutdown period.
  • Wet expanders’ u-tubes have been reinstalled; cold compressors will be re-installed as necessary.
  • Reciprocating cold compressors are being overhauled for A4, B1, and C4.
  • Tevatron Cryogenic Transfer Line between D1 and D2 was leak checked due to vacuum rise last week and no leaks were found. We will monitor vacuum over time.
  • D2 EV101 main bldg shutoff valve leak-thru was investigated and actuator alignment fixed it. B4EV101 valves leaks thru very badly and needs to be investigated; repairs may necessitate isolating whole sector helium discharge header.
  • All helium to atm leaks on safety leads and conning towers due to fittings were fixed, except for A1, A3, and F4 .

Facility Engineering Services

  • Zebra mussel cleaning began on May 8.

Instrumentation

Work continues.

RF

Work continues.

Controls

Work continues.

EE Support

Work continues

Alignment

  • Tevatron
    • Scheduled A 48 Collimator
    • Completed Add on F17-3 Kicker as found.
  • Main Injector
    • Completed
    • Scheduled Add on OTR P150 line
  • Recycler
    • Completed
    • Scheduled Add on Job for Monday
  • 8Gev
    • Completed
    • Scheduled Add On Level run on floor point.
  • P-Bar
    • Completed deep rod installation
  • MiniBooNE
    • Completed
  • NuMI
    • Completed
  • Booster
    • Completed
  • 400Mev
    • Completed
  • Switchyard
    • Scheduled Add on Check the alignment of elements in Switchyard manhole enclosure F3.
  • D0
    • Closing Stick Mic.
    • VSTAR survey of detector (center beam) position
    • Level-Run vertical measurement of select Collision Hall
  • Add Ons
    • D0 - Make Level-Run vertical measurement of select Collision Hall monuments West Wall
    • Main Injector Align barrier cavity at MI604 location. Completed
    • Fess - the Batavia Rd. Guardhouse project. Completed
    • Fess - MI Ponds C&D project will replace existing manholes and pipe.
    • Switchyard - Check the alignment of elements in Switchyard manhole enclosure F3.
    • CDF Scheduled CDF Low Beta Alignment
    • 8Gev Magnet re-alignment Moved Completed
    • 8Gev Elevation run (Wednesday) Completed
    • NuMI Target Hall check for deformation

Schedule for Next Week

 

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 05/15/2006    email Fermilab